# Enhance Mobile Agents Thinking Process Via Iterative Preference Learning

### v2

@@ 4.1 Experiments Setups
Dataset. AITZ (zhang2024android, ) is a high-quality trajectory set filtered and re-annotated from AITW (rawles2024androidinthewild, ), containing four subsets: general, install, single, and Google apps, which also includes five types of actions: click, input, scroll, press, and stop. AMEX (chai2024amex, ) uses the same apps and action space as AITZ, but its task instructions are more complex and detailed, with an average trajectory length of 15+. AndroidControl (li2024effects, ) includes different task types compared to the previous two datasets. In addition, it contains extra actions such as wait and completely unseen out-of-domain test splits.
Evaluation. For evaluation, we use Step.Acc as metrics, consistent with Auto-GUI(zhang2023you, ), measures the agent’s performance and uses Action Type to assess the degree of action type matching. This metric effectively evaluates the model’s planning ability.
Baselines. Following prior work(wu2024atlas, )(qin2025ui, ), we use Qwen2-VL-7B (wang2024qwen2, ) as the backbone of our model. We select CogAgent (hong2024cogagent, ), AUTO-GUI, Shpagent, OS-Atlas, UGround, UI-Tars and FedMobileAgent as baseline agents. GUI continuous pre-training agents can be further divided into two categories: (1) training the model as a GUI grounding agent, such as OS-Atlas-7B. (2) training the model as a general GUI agent, such as UI-Tars. More details are provided in Appendix D.
@@ 4.2 Main result
AITZ. As shown in Table 1, MobileIPL achieves SoTA performance on most metrics. The reason for the lower PRESS Acc. is discussed in Section F and Appendix H. Multiple rounds of T-DPO improve MobileIPL by more than 10 points (55.40% -> 69.15%) compared to the seed model MobileIPL and ablation model Qwen2-VL-7B (60.36% -> 69.15%). Compared to continuous pre-training agents such as Falcon-UI, which is pre-trained on three million GUIs, MobileIPL still surpasses a performance difference of 0.05%. The amount of training data required by our method is substantially smaller than that used by these pre-training approaches. The GUI pre-training Grounding model UGround underperforms its base model Qwen2-VL on AITZ downstream tasks (60.19% < 60.36%), but UI-Tars performs better (65.61% > 60.36%) because of its more varied pre-training tasks.
@@
AMEX. As shown in Table 2, except for the SHEIN and ToDo applications, our agent achieves SoTA performance on all other applications, surpassing the previous SoTA pre-training SphAgent by 3.58% (70.71% -> 74.29%). This demonstrates the effectiveness of our method. The CoaT reasoning process significantly improves Qwen2-VL performance (69.01% -> 70.93%), which is nearly on par with the GUI domain pretraining model SphAgent (70.71% $vs.$ 70.93%). The improvements brought by instruction evolution and IPL training are well-balanced. Among the 9 evaluated apps, MobileIPL achieves either the best or second-best performance on 7 of them, while the second-best agent, Sphagent, exhibits relatively unbalanced performance (exhibits high performance on SHEIN and ToDo).
@@
Table 4: High-level instruction results on AndroidControl in-domain and unseen OOD subsets.
@@
AndroidControl. As shown in Table 4. MobileIPL achieves SoTA in Step.Acc (72.7%), surpassing continual pre-training agents of the GUI domain such as OS-Atlas (+1.5%) and UI-Tars (+0.2%). However, since these models gain Grounding capabilities during the pretraining phase, MobileIPL still underperforms in element grounding compared to OS-Atlas (-1.5%) and UI-Tars (-3.5%). As shown in Table 4, in unseen OOD, the pre-trained model OS-Atlas shows only limited improvements over Qwen2-VL (+2.1%, -0.7%, +2.1%), whereas MobileIPL achieves significant gains (+3.5%, +8.4%, +8.1%). Additionally, compared to in-domain data, MobileIPL exhibits less performance degradation in the out-domain setting. For example, OS-Atlas experiences a 10.5% drop in the Unseen app subset (71.2% -> 60.7%), while MobileIPL only decreases by 3.6% (73.6% -> 70.0%).
@@ 4.3 Ablation Study
Since our method incorporates Instruction Evolution (Evo), we conduct an ablation study to analyze the contribution of each component. As shown in Table 5, removing IPL leads to a significant drop in overall performance from 69.2% to 60.4%, demonstrating that IPL plays a crucial role. In contrast, removing Evo results in only a marginal performance decrease (-0.1%), indicating that Evo does not directly enhance the model’s performance. We also examine the effects of dialogue (Dia) and history (His) information.
@@
Standard Self Training. Additionally, we remove negative samples from IPL-R1, training the model using only fully correct samples as standard self-training does. This results in a 4.0% performance drop, suggesting that negative samples help the model learn how to reason rather than merely memorize (SFT). Furthermore, compared to training on the entire trajectory with naive DPO, the model’s performance drops from 65.4% to 60.3%, which proves the effectiveness of the CoaT-tree sampling and thinking process optimization.
Low Resource. We also perform low-resource on AITZ, sampling 1/2 and 1/5 of the training data. As shown in table 5, using only half of the data, the first round of IPL training already outperformed the best results achieved by the original CoAT-SFT and naive DPO training. Furthermore, when using only one-fifth of the data, the second round of IPL training surpassed the performance of CoAT-SFT (- IPL,Evo), demonstrating the effectiveness of our method even in low-resource scenarios.
@@ 4.4 Discussion and Analysis
Reasoning Space Sampling. To evaluate the instruction evolution, we analyze the diversity of the sampling space for Random 1000 steps, the standard deviation of encoded embeddings, the dimensionality-reduced distribution, and the distribution of $S^{(i)}$ mentioned in Section 3.3. As shown in Figure 3 (a), the thoughts after instruction evolution exhibit a broader space than direct SFT. Additionally, the embedding standard deviation within each tree increases significantly compared to the original data (+ 0.158). The diversified outputs do not negatively impact the agent’s reasoning process, while the proportion of action sampling that includes the correct answer improves from 72.7% to 77.9%. The bottom-right subplot reflects the distribution of output accuracy. Consistently Correct indicates that all samples for the current step match the golden answer, while Consistently Error is the opposite. Both represents cases where some samples are correct while others are incorrect, which serves as an ideal source for constructing DPO pairs. Compared to 47% on the evolved data, the agent achieves 68.7% convergence on the original data but exhibits a strong polarization(4%). Three-stage instruction evolution significantly expands the sampling space (from 4% to 31%), proving that it simultaneously improves both the diversity and quality of reasoning. More details are in Appendix G.
Figure 3: (a) Reasoning diversity before and after instruction evolution (left top) and the distribution of standard deviation and accuracy (left bottom). (b) Step.Acc changes for action types in AITZ across IPL iterations (right top). (c) The performance of UI-Tars-7B and Qwen2-VL-7B on AITZ as the seed model with multi-round IPL training (right bottom).
@@
Parameters Searching. We conducted an ablation study on the impact of the sampling number, and based on the trade-off between performance and computational cost, we chose 3 as the final setting. As shown in the table 6, increasing the number of samples generally leads to better model performance. However, since our framework adopts a tree structure, increasing the sampling number from 3 to 4 causes the minimum number of tree nodes to grow significantly from $273^{3}=$ 27 to $644^{3}=$ 64. Despite this sharp increase, the performance improvement is limited (less than 1%). Therefore, we adopt a sampling number of 3 for the final experiments.
IPL Scaling. Although the total Step.Acc increases with IPL iterations, not all action types follow this upward trend. As shown in Figure 3 (b), from the seed model to the first round of IPL, the accuracy of PRESS drops significantly (58.22% -> 23.49%), while CLICK increases (53.26% -> 71.12%). Inspired by action equivalence matching in AndroidControl, we attempt to analyze the reasons behind this phenomenon. In fact, multiple actions on the same GUI can result in an equivalent outcome. For example, after entering a search query, PRESS ENTER and directly click on a dropdown suggestion with the same keyword both navigate to the same page. We analyze the proportion of steps where the ground truth action is PRESS, but the model outputs a different action. We find that when PRESS accuracy drops to 23.49%, CLICK accounts for 85.10% of the errors in the lower part of the figure. As more positive PRESS pairs (14.5% -> 18.7%) and less CLICK data (43.3% -> 37.4%) are sampled in subsequent IPL iterations, this error proportion gradually decreases to 73.7%, and PRESS Acc. increases to 47.78%. For detailed statistical data, please refer to Appendix F.
@@ 5 Conclusion
In this paper, we propose Mobile Iterative Preference Learning (MobileIPL), a self-training GUI agent framework that incorporates instruction evolution, iterative sampling in the CoaT-tree, and a rule-based reward. We extensively evaluate MobileIPL on the AITZ, AMEX, and AndroidControl benchmarks, demonstrating its effectiveness. Furthermore, MobileIPL exhibits strong generalization capabilities on the unseen subsets of AndroidControl. Experiments show that instruction evolution does not directly improve downstream task performance, but it effectively enhances the diversity of the model’s outputs. Finally, Continuous Pre-training experiments confirm the mutual reinforcement between MobileIPL and pre-training, leading to enhanced performance.
@@ Appendix B Selection of Seed Policy Model
$\small Acc_{S}=\frac{\sum_{i=1}^{|\mathcal{T}|}\Big{|}\big{\{}e_{j}^{(i)}\mid a^{(i)}\sim e_{j}^{(i)},e_{j}^{(i)}\in\mathcal{S}^{(i)}\big{\}}\Big{|}}{\sum_{i=1}^{|\mathcal{T}|}|\mathcal{S}^{(i)}|}$ (14)
@@
$\small Dev_{{S}^{(i)}}=\frac{1}{T}\sum_{t=1}^{T}\mathrm{StdDev}\left(\mathbf{E}(\hat{s}^{(k)}_{t})\mid k=1,\dots,\mathcal{K}\right)$ (15)
@@ Appendix C Rule-based Reward Design
Strong Reward: We follow the ReFT [4] score to define strong reward signals, assigning values of 1 and 0, corresponding to fully correct and completely incorrect reasoning paths, respectively. In ReFT, a supervision signal of 0.1 encourages the model to produce a final answer following the predefined format. In our approach, this signal is repurposed to reward action type matching. Meanwhile, an additional $v_{format}$ reward is introduced to encourage proper formatting of input and click actions.
@@
Figure 4: The proportion of CLICK and PRESS actions selected as positive data during the first four rounds of IPL.
@@ Appendix D Experiment Setup
CoaT Multi-round Dialogue Prompts.
@@ Appendix E Human Annotation
Appendix F IPL Scaling Analysis
In the first four rounds of DPO pairs, the number of CLICK positive samples is 1731, 1496, 1485, 1475, while the number of PRESS positive samples is 580, 749, 579, 315. During the first round of IPL, when PRESS accuracy dropped to 10.18%, CLICK has the highest count at 1731. In the following round, CLICK decreased by 235 (1731 → 1496), while PRESS increased by 169 (580 → 749). After this round of training, PRESS Acc. recovers to 31.07%, leading to a gradual decline in the proportion of this action type in the following round of sampling. In subsequent rounds, their ratio gradually stabilizes. For detailed sampling data changes, please refer to Figure 4.
Appendix G Iterative Sampling in the CoaT-tree
@@ Appendix H Case Study
Appendix I Broader Impact
This paper presents an iterative preference learning framework that includes iterative sampling, instruction evolution, and rule-based reward design. The proposed method constructs a CoaT-tree through iterative sampling and leverages rule-based reward with tree-based credit assignment to generate think-level DPO pairs for training, aiming to improve the model’s reasoning process.To mitigate overfitting during the warm-up supervised fine-tuning (SFT) stage, we further introduce a three-stage instruction evolution strategy to enhance the diversity of model outputs.Extensive experiments on three benchmark datasets—AITZ, AMEX, and AndroidControl—demonstrate that our method outperforms existing state-of-the-art models, including those based on continual pretraining. From a broader perspective, this work contributes to the development of more robust and interpretable reasoning in Mobile GUI agents, which may benefit downstream applications such as task automation, mobile accessibility tools, and human-computer interaction. By explicitly optimizing intermediate reasoning steps, the framework encourages more transparent and verifiable agent behavior.
Appendix J Limitations
Our iterative preference learning method involves multiple rounds of sampling the entire training data, which takes much longer compared to direct SFT, especially when the amount of training data is very large. In our experiments, since some of the continuous pre-training models in the GUI domain do not provide the format and prompt of the training data, the reproduction results may be biased.
Appendix K Ethics Statement

### v3

@@ 4.1 Experiments Setups
Dataset. AITZ (Zhang et al., 2024b) is a high-quality trajectory set filtered and re-annotated from AITW (Rawles et al., 2023), containing four subsets , which also includes five types of actions AMEX (Chai et al., 2024) uses the same apps and action space as AITZ, but its task instructions are more complex and detailed, with an average trajectory length of 15+. AndroidControl (Li et al., 2024a) includes OOD datasets, such as app unseen and task unseen.
Metrics. For evaluation, we use Step.Acc as metrics, consistent with Auto-GUI(Zhang & Zhang, 2023), measures the agent’s performance and uses Action Type to assess the degree of action type matching. This metric effectively evaluates the model’s planning ability.
Baselines. Following prior work(Wu et al., 2024)(Qin et al., 2025), we use Qwen2-VL-7B (Wang et al., 2024b) as the backbone of our model. We select CogAgent (Hong et al., 2024), AUTO-GUI, Shpagent, OS-Atlas, UGround, UI-Tars and FedMobileAgent as baseline agents. GUI continuous pre-training agents can be further divided into two categories: (1) training the model as a GUI grounding agent, such as OS-Atlas-7B. (2) training the model as a general GUI agent, such as UI-Tars. More details are provided in Appendix D.
@@ 4.2 Main result
AITZ. As shown in Table 1, MobileIPL achieves SoTA performance on most metrics. The reason for the lower PRESS Acc. is discussed in Section 6 and Appendix H. Multiple rounds of T-DPO improve MobileIPL by more than 10% (55.40% -> 69.15%) compared to the seed model MobileIPL and Qwen2-VL-7B (60.36% -> 69.15%). Compared to continuous pre-training agents such as Falcon-UI, which is pre-trained on three million GUIs, MobileIPL still surpasses a performance difference of 0.05%. The amount of training data required by our method is substantially smaller than that used by these pre-training approaches.
@@
AMEX. As shown in Table 2, MobileIPL surpasses the previous SOTA model, SphAgent-7B, by 3.58%. It also outperforms the baseline model (Qwen2-vl+CoaT) by 3.36%. Additionally, MobileIPL surpasses OS-Atlas (+3.69%) and UI-Tars (+3.69%), both of which also use Qwen2-vl as the backbone. With the incorporation of CoaT, the baseline model Qwen2-vl shows an increase of 1.92%, demonstrating the effectiveness of CoaT patterns. In summary, these results confirm that MobileIPL delivers significant improvements over existing models in long trajectory scenarios.
@@
Table 4: High-level instruction results on AndroidControl in-domain and OOD subsets.
@@
AndroidControl. As shown in Table 4, MobileIPL achieves SOTA performance in Step.Acc (72.7%), reaching the SOTA model Falcon-UI with fewer data. MobileIPL also outperforms continual pre-training agents in the GUI domain, such as OS-Atlas (+1.5%) and UI-Tars (+0.2%). Compared to the baseline model Qwen2-VL(SFT), MobileIPL not only improves Mobile Agent performance but also enhances grounding by 8.5%. As shown in Table 4, MobileIPL continues to achieve SOTA performance in unseen OOD settings, demonstrating strong generalization. In contrast, compared to performance in the IDD domain, the pre-trained model OS-Atlas shows a significant drop. MobileIPL exhibits less performance degradation in out-of-domain settings. We also ran GRPO with Qwen2-VL under the same computational resources, and found OOD performance similar to MobileIPL, because both are self-training. However, MobileIPL still outperforms GRPO in all subsets.
@@ 4.3 Ablation Study
To test the effectiveness of IPL and instruction evolution, we conducted ablation experiments. First, removing IPL and using only SFT caused performance to drop from 65.4% to 60.4%, compared to the first round of MobileIPL, highlighting the crucial role that IPL plays. Next, removing instruction evolution led to a 2.5% drop in IPL performance in the first round. This occurs because, without evolution, the model generates fewer training samples (156,418 -> 113,239). And as shown in Figure 3 (a), without instruction evolution, the diversity of model outputs decreased, causing a drop in IPL performance. This further confirms that instruction evolution is crucial for improving IPL.
@@
Standard Self Training. Additionally, we remove negative samples from IPL-R1, training the model using only fully correct samples. This results in a 4.0% performance drop, suggesting that negative samples help the model learn how to reason rather than merely memorize (SFT). Furthermore, training on the entire trajectory with navie DPO reduces performance from 65.4% to 60.3%. Compared with SFT trained on CoaT tree positive data (–IPL Negative), naive DPO is still 1.1 % lower, confirming the effectiveness of CoaT-tree sampling and the thinking-process optimization.
Low Resource. We also perform low-resource on AITZ, sampling 1/2 and 1/5 of the training data. As shown in table 5, using only half of the data, the first round of IPL training already outperformed the best results achieved by the original CoAT-SFT (-IPL) and naive DPO training. Furthermore, when using only one-fifth of the data, the second round of IPL training surpassed the performance of CoAT-SFT (-IPL), demonstrating the effectiveness of our method even in low-resource scenarios.
@@ 4.4 Discussion and Analysis
Reasoning Space Sampling. To evaluate the instruction evolution, we analyze the diversity of the sampling space for Random 1000 steps, the standard deviation of encoded embeddings, the dimensionality-reduced distribution, and the distribution of $S^{(i)}$ mentioned in Section 3.3. As shown in Figure 3 (a), the thoughts after instruction evolution exhibit a broader space than direct SFT. Additionally, the embedding standard deviation within each tree increases significantly compared to the original data (+ 0.158). The diversified outputs do not negatively impact the agent’s reasoning process, while the proportion of action sampling that includes the correct answer improves from 72.7% to 77.9%. The bottom-right subplot reflects the distribution of output accuracy. Consistently Correct indicates that all samples for the current step match the golden answer, while Consistently Error is the opposite. Both represents cases where some samples are correct while others are incorrect, which serves as an ideal source for constructing T-DPO pairs. Compared to 47% on the evolved data, the agent achieves 68.7% convergence on the original data but exhibits a strong polarization(4%). Three-stage instruction evolution significantly expands the sampling space (from 4% to 31%), proving that it simultaneously improves both the diversity and quality of reasoning. More details are in Appendix F.
Figure 3: (a) Reasoning diversity before and after instruction evolution (left top) and the distribution of standard deviation and accuracy (left bottom). (b) Step.Acc changes for action types in AITZ across IPL iterations (right top). (c) The performance of UI-Tars-7B and Qwen2-VL-7B on AITZ as the seed model with 4-round IPL training (right bottom).
@@
Parameters Searching. We conducted an ablation study on the impact of the sampling number ($\mathcal{K}$) per stage and iterative round number ($R$). As shown in the table 6, increasing the number of samples generally leads to better model performance. However, since our framework adopts a tree structure, increasing the sampling number from 3 to 4 causes the minimum number of tree nodes to grow significantly from $3^{3}=27$ to $4^{3}=64$. Despite this sharp increase, the performance improvement is limited (less than 1%). Therefore, we adopt a sampling number of 3 for the final experiments. Regarding the number of rounds, we observe that both IPL performance and the size of the self-training dataset converge after several iterations. We therefore select the convergence round as our default setting. Additional details on computational cost are provided in Appendix G.
IPL Scaling. Although overall Step.Acc increases across IPL iterations, not all action types follow this trend. As shown in Figure 3(b), from the seed model to the first IPL round, PRESS accuracy drops sharply (58.22% → 23.49%), whereas CLICK rises (53.26% → 71.12%). In the second round, however, PRESS accuracy rebounds. This stems from the severe underrepresentation of PRESS actions early on: the proportion of PRESS samples in the preference data grows from 1.6% (1 round) to 10.9% (2 round) as training progresses. With greater reasoning diversity and more PRESS-related examples, the model gradually learns PRESS behaviors and recovers accuracy in later rounds.
@@ 5 Conclusion
In this paper, we propose Mobile Iterative Preference Learning (MobileIPL), a self-training GUI agent framework that incorporates instruction evolution, iterative sampling in the CoaT-tree, and a rule-based reward. We extensively evaluate MobileIPL on the AITZ, AMEX, and AndroidControl benchmarks, demonstrating its effectiveness. Furthermore, MobileIPL exhibits strong generalization capabilities on the OOD subsets of AndroidControl. Experiments show that instruction evolution increases output diversity, generates more training data in IPL, and thereby improves IPL performance. Finally, Continuous Pre-training experiments confirm the mutual reinforcement between MobileIPL and pre-training, leading to enhanced performance.
6 Ethics Statement
7 Reproducibility statement
All models and datasets used in this paper are open-source. The full experimental setup is detailed in Appendix D. Unless noted, all experiments use the same settings. We describe compute resources in Appendix G. Overall, these practices make our results reproducible.
@@ Appendix B Selection of Seed Policy Model
$Acc_{S}=\frac{\sum_{i=1}^{|\mathcal{T}|}\Big|\big\{e_{j}^{(i)}\mid a^{(i)}\sim e_{j}^{(i)},e_{j}^{(i)}\in\mathcal{S}^{(i)}\big\}\Big|}{\sum_{i=1}^{|\mathcal{T}|}|\mathcal{S}^{(i)}|}$ (15)
@@
$Dev_{{S}^{(i)}}=\frac{1}{T}\sum_{t=1}^{T}\mathrm{StdDev}\left(\mathbf{E}(\hat{s}^{(k)}_{t})\mid k=1,\dots,\mathcal{K}\right)$ (16)
@@ Appendix C Rule-based Reward Design
Strong Reward: We follow the ReFT (Luong et al., 2024) score to define strong reward signals, assigning values of 1 and 0, corresponding to fully correct and completely incorrect reasoning paths, respectively. In ReFT, a supervision signal of 0.1 encourages the model to produce a final answer following the predefined format. In our approach, this signal is repurposed to reward action type matching. Meanwhile, an additional $v_{format}$ reward is introduced to encourage proper formatting of actions.
@@
Text F1 vs. Text-Embedding Similarity: We replaced the F1-based reward with a BERT-based semantic reward and evaluated both variants. As shown in Table 8, the F1 reward outperforms the BERT embedding reward across all metrics, with the largest gain on TYPE ACTION (acc) (+4.53%). This aligns with the importance of exact keyword matching in GUI input, indicating that F1 is better suited than semantic similarity for reward design in mobile UI input scenarios.
Table 8: Text F1 vs. text-embedding similarity: ablation on reward design. Best results are in bold.
Reward Type TYPE (type) TYPE (acc) TOTAL (type) TOTAL (acc)
BERT 84.17 76.70 77.81 62.49
F1 87.78 81.23 78.74 65.37
@@ Appendix D Experiment Setup
CoaT Multi-turn Dialogue Prompts.
@@ Appendix E Human Annotation
Appendix F Iterative Sampling in the CoaT-tree
@@
Appendix G Computational Cost of IPL Across Iterations
Table 9: Data size and training time across iterative rounds.
Standard SFT Round 1 Round 2 Round 3 Round 4 Round 5
Data Samples 63,158 156,418 34,010 28,491 17,027 17,780
Time (hours) 2.8 7.9 2.0 1.7 1.0 1.0
As shown in Table 9, while the first IPL iteration incurs higher cost due to the larger volume of preference data, subsequent iterations are significantly lighter. The training time per round decreases rapidly, as the model generates fewer low-quality samples and requires fewer updates. In fact, by the third iteration, the training time becomes comparable to or even lower than the initial supervised fine-tuning (SFT) stage.
Therefore, the overall compute overhead of IPL remains moderate and manageable, especially considering its performance gains. Compared to SFT, IPL introduces only a modest increase in compute, but brings substantial improvements in reasoning and generalization.
@@ Appendix H Case Study
Appendix I Usage of LLM statement
This paper utilized an LLM to improve the clarity and fluency of the text.